When 23.8 g of a group 1 carbonate was heated strongly, it decomposed to form the oxide of the group 1 element (call it element X) and 2.70 L of carbon dioxide gas, collected at 25*C. The density of carbon dioxide gas at that temperature was 1.20 g/L. Determine the quantity of the metal , X, and write the balanced equation for the reaction.
1
Expert's answer
2019-10-02T08:02:49-0400

The equation of decomposition is:

"X_2CO_3\to X_2O+CO_2"

Mass of "CO_2" formed"=1.20\frac{g}{L}\times2.70L=3.24g"

Moles of "CO_2" formed"=\frac{3.24g}{44\frac{g}{mol}}=0.0736mol"

Molar mass of "X_2CO_3=\frac{23.8g}{0.0736mol}=323.369\frac{g}{mol}"

So atomic mass of "X=\frac{323.369-60}{2}=131.6845amu"

Hence quantity of metal "X=131.6845\times0.0736=9.69g"